OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2)

OpenStudy (anonymous):

so the slope of the line can be found with -b/2a so the slope is -(-3)/6 or 1/2 parallel lines have the same slope so now that we have the slope, we can use point slope form y-y1=m(x-x1) y+2=1/2(x-4) y+2=1/2x-2 and i guess you want this in slope intercept form so y=1/2x-4

OpenStudy (anonymous):

explain -b/2a mechanism

OpenStudy (anonymous):

sorry its supposed to be -a/b so the slope is -3/-1=3 the same other reasoning would follow jsut change the slop -b/2a is to find the axis of symmetry for a parabola sorry! y-y1=m(x-x1) y+2=3(x-4) y+2=3x-12 y=3x-14

OpenStudy (anonymous):

whats a and what is b a is the slope x, b is the value with y ?

OpenStudy (anonymous):

standard form is ax+by=c a is the coefficient to x b is the coefficient to y

OpenStudy (anonymous):

yes thanks I understand now

OpenStudy (anonymous):

sorry about the mix up

OpenStudy (anonymous):

not a problem for me, I have not applied any of it yet ;)

OpenStudy (anonymous):

still really sorry i think i was helping someone with parabolas earlier and it was stuck in my head